OpenStudy (truehawaiian):

Can someone help me... Evaluate the expression a=6 and b=39. b-5a

OpenStudy (anonymous):

b-5a=39-5*6 do you want to continue??

OpenStudy (truehawaiian):

Yes please...

OpenStudy (anonymous):

39-30=9

OpenStudy (truehawaiian):

Ah... Thank you! I didn't know what I was suppose to do from there my teacher just left us with that... Thank you though I know what to do from here on thank you...

OpenStudy (anonymous):

you're welcome